OpenStudy (anonymous):

What is the range of the graph of y = 3(x – 3)2 + 4? y less than or equal to 4 y greater than or equal to 4 y greater than or equal to 3 y less than or equal to 3

OpenStudy (anonymous):

I've never learned anything about the range of a graph :C

OpenStudy (anonymous):

well read the question again

OpenStudy (anonymous):

i think there is a mistake

OpenStudy (anonymous):

I don't know if maybe it's talking about where y starts and continues on the graph? Which in that case, shouldn't it be y is greater than or equal to 4?

OpenStudy (anonymous):

well there is no limit of y as u can see it's continous

OpenStudy (anonymous):

Which is why I think it may be the second choice.

OpenStudy (anonymous):

3(x-3)2=6(x-3) ram i reading it corectly

OpenStudy (anonymous):

y=3(x-3)^2+4

OpenStudy (anonymous):

is it ^-2

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

y greater than or equal to 4

OpenStudy (anonymous):

Thank you, that's what I thought!

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

no i don't think how ru saying

OpenStudy (anonymous):

i mean why do u think

OpenStudy (anonymous):

When I graphed it, it was showing how it's continuous. So it can't be any choice with less than in it since the parabola is facing up. And it can't be any choice with 3 either, because the vertex starts at (3,4).